# An analytical exploration of the optomechanical attractor diagram and of limit cycles

### v1

@@ III Limit cycles
where $\mu(r)$ and $D(r)$ are the drift and diffusion coefficients. It is worth mentioning that, while in the main text in [12] the case of the $Q$-function was considered, the more general parametrized phase-space quasiprobability distribution was presented as well and the drift coefficient mantains the same form regardless of the initial quasiprobability distribution used to describe the system’s density operator. However, the diffusion coefficient is specific to the phase-space distribution [12], but it is equally amenable to our Bessel summation formulas, as we show below for the case of the Wigner function.
@@ III.1 Drift and diffusion coefficients of the effective equation of motion
We start by providing an exact calculation of the infinite sums describing the drift and diffusion coefficients. They are given by [12]
@@
Here $r$ represents the amplitude, $E=\hbar\omega_{m}r^{2}$ is the energy, $g_{0}$ is the single-photon coupling strength and the dimensionless coupling $\eta=2g_{0}/\omega_{m}$ is also used to characterize the optomechanical coupling [12], $\kappa$ is the cavity amplitude decay rate and $\Delta=\omega_{L}-\omega_{c}$ is the detuning between the laser frequency $\omega_{L}$ and the bare cavity resonance $\omega_{c}$. The mechanical damping rate parameter is denoted by $\gamma$.
The term effective detuning $\Delta_{\rm eff}$ represents the fact that there is a dynamical dependence of the effective detuning on the intracavity amplitude and a corresponding nonlinear dependence on the cycle amplitude. In particular:
$\Delta_{eff}=\Delta+2K\sum_{n}\left|\alpha_{n}\right|^{2},$
where $K$ denotes the strength of an effective Kerr nonlinearity with $K=\frac{g_{0}^{2}}{\omega_{m}}$. Indeed, in the formalism of [12] the cavity acquires an effective Kerr nonlinearity that is encoded in $\Delta_{eff}$. Because the Kerr nonlinearity induces a different effective detuning for the mean field $\alpha$ than for the fluctuations, we also have:
$\tilde{\Delta}_{eff}=\Delta+4K\sum_{n}\left|\alpha_{n}\right|^{2}.$
The static approximation for the detuning is valid for a negligible Kerr parameter $K$ or in the weak driving limit. Indeed, in the $\left|\alpha(t)\right|^{2}\ll 1$ limit the bare detuning is just shifted by the constant Kerr term and there is no need to distinguish between $\Delta_{eff}$ and $\widetilde{\Delta}_{eff}$. The shift $\Delta_{eff}=\Delta+K$ appears naturally in the framework of [12] (see also [16, 25]) and explains the appearance of limit cycles for specific particular values given in [12] and also, for example, for limit cycles on the blue sideband in [25]. So, already in the static approximation model, this classical Kerr shift is of importance.
@@
A discussion of both the static and the dynamical setting is alternatively presented throughout the whole discussion in the paper [12]. Here, we first discuss the static case, valid for $\left|\alpha(t)\right|^{2}\ll 1$ and then in Section B.1 below, we discuss the dynamical case by evaluating $\Delta_{\rm eff}(r)$ in full, in contrast to the sideband-resolved approximation of [12]. Indeed, these sums have been evaluated in a close-to-resonance approximation [12], where you discard an infinite number of terms. Using the summation formulas of appendix A, we can now provide exact expressions for the complete resummation. For the drift, we obtain
@@
Details of the calculation are given in appendix A and below we will discuss its asymptotic value for large $\eta r$. For the diffusion coefficient, we find
@@
$D_{a}(r)=-\frac{\pi}{2\omega_{m}}\,g_{0}^{2}E^{2}{\rm Im}\left[B\left(\frac{J_{\nu}(\eta r)J_{-\nu}(\eta r)}{\sin(\pi\nu)}+\frac{J_{1+\tilde{\nu}^{*}}(\eta r)J_{-1-\tilde{\nu}^{*}}(\eta r)}{\sin(\pi\tilde{\nu}^{*})}\right)\right]$ (28)
@@
$\tilde{\nu}=\frac{\tilde{\Delta}_{\rm eff}}{\omega_{m}}+\frac{i\kappa}{\omega_{m}}$ and the constants $B$, $C_{1}$, $C_{2}$, $C_{3}$ are given in appendix A.
@@ III.1.1 Diffusion in the Wigner case
The Fokker-Planck equation (21) can also be written for the Wigner function. It has the same drift coefficient, and a diffusion given by (see (B.20) in [12])
@@
All the arguments of the Bessel functions are $\eta r$. This expression plays a relevant role in the dynamics of limit cycles. The infinite sums can again be computed exactly by using the summation formulas of the appendix. The final result is a long expression and we will not need it here, but $D_{W}$ drastically simplifies when $\tilde{\Delta}_{\rm eff}=\Delta_{\rm eff}$. This identification is justified in the sideband-resolved regime studied in [12]. We obtain (see Appendix for details)
@@
In fig 2 we compare our exact formula for $D_{W}$ with $D_{W}^{(0)}$ for $\kappa=0.1$, $\omega_{m}$ (resolved sideband) and two different values of $\Delta_{\rm eff}$. We see that the approximation (32) is very inaccurate for $\Delta_{\rm eff}=0.6\,\omega_{m}$ A closer examination shows that the approximation (32) is no longer applicable (with an error higher than 20 per cent) for $\Delta_{\rm eff}>0.4\,\omega_{m}$.
@@
Figure 2: Optical part of the Wigner diffusion from approximation (32) (blue) and for the exact formula (31) (red) for $\kappa=0.1\omega_{m}$. a) $\Delta_{\rm eff}=0.1\,\omega_{m}$. The approximation (32) reproduces with high accuracy the exact curve. b) $\Delta_{\rm eff}=0.6\,\omega_{m}$. The approximation (32) is no longer applicable as higher $n$ terms become important.
III.2 Asymptotic formulas, classical optomechanics cycles and the dynamical setting $\Delta\equiv\Delta_{\rm eff}(r)$
@@
To simplify the notation, here we wrote $\Delta\equiv\Delta_{\rm eff}(r)$. This formula exhibits the relevant periods of oscillations in $r$ and in $\Delta/\omega_{m}$. A surprising effect predicted by the asymptotic formula for $\mu(r)$ is that the second term in (34) identically vanish for
@@
Thus, for these special values of $\Delta$, $\mu(r)=-\gamma r+O(1/r)$. Oscillations are suppressed by a factor $1/r$ originating from subleading terms in the asymptotic expansion of the Bessel functions. This effect is shown in fig. 3 and is a result of the full resummation. It is not seen if one truncates the sum to few values of $n$, as in a close-to-resonance approximation. It would be interesting to see this effect in experiments.
@@ III.2.2 Dynamical setting
The infinite sum can again be computed using the summation formulas of the appendix by noting that this can be re-writen as
@@
This can be solved by assuming that $\Delta_{\rm eff}\ll\omega_{m}$. Then
@@
This formula encodes the information of all secondary resonances.
@@ IV Outlook
The technical advance of this paper lies in the formulation of a compact, simple formula that comprehensively encompasses all terms within the summation. This provides an analytical understanding of the parameter space and offers insight into the underlying physics of the various processes. Additionally, it facilitates the application of asymptotic expressions, resulting in expressions that are expressed in terms of uniquely elementary functions. Since these are simple trigonometric functions, further elucidation, for example of the zeros [10], becomes a particularly straightforward task, helping to identify special points in regions of the parameter space in limits where a resolved sideband approximation, which retains only a few resonant terms in the summation, does not reach. There are also open problems already in the context of the present discussion. For example, whether the more complete analysis of Wigner diffusion presented here also allows a more detailed analysis of energy fluctuations by evaluating the Fano factor, as in [12], but with full summation. All these considerations require no further mathematical developments than those presented here or in the mathematically related works [9, 10].
@@ Acknowledgements.
We thank Professors David Pérez-García, Oriol Romero-Isart and Patrick Maurer for discussions and correspondence at different stages of this work. The work of MT was started at Universidad Complutense de Madrid (UCM) and continued and finished after moving to the Shanghai Institute for Mathematics and Interdisciplinary Sciences (SIMIS). JGR acknowledges financial support from grant 2021-SGR-249 (Generalitat de Catalunya) and by the Spanish MCIN/AEI/10.13039/501100011033 grant PID2022-126224NB-C21. MT acknowledges previous financial support from FEI-EU-22-06, funded by Universidad Complutense de Madrid, and grant PID2020-113523GB-I00, funded by the Spanish Ministry of Science and Innovation.
Appendix A Bessel summations and diffusion coefficients

### v2

@@ III Limit cycles
where $\mu(r)$ and $D(r)$ are the drift and diffusion coefficients. All this is explained in more detail in Appendix A, summarizing the basic formalism developed in [12]. It is worth mentioning that, while in the main text in [12] the case of the $Q$-function was considered, the more general parametrized phase-space quasiprobability distribution was presented as well and the drift coefficient mantains the same form regardless of the initial quasiprobability distribution used to describe the system’s density operator. However, the expression for the diffusion coefficient depends on the phase-space distribution. This can be seen in the expression given for the case when the phase-space distribution is the Wigner function [12]. This more complex expression for the diffusion coefficient can also be fully evaluated using Bessel summation formulas, as we demonstrate explicitly below for diffusion corresponding to the Wigner function.
@@ III.1 Drift and diffusion coefficients of the effective equation of motion
We start by providing an exact calculation of the infinite sums describing the drift and diffusion coefficients in the Fokker-Planck formulation of [12]. We explain the origin, providing also context for these expressions in Appendix A (where the notation and conventions are also further discussed). They are given by [12]
@@
Here $r$ represents the amplitude, $E$ is the laser drive rate, $g_{0}$ is the single-photon coupling strength, the dimensionless coupling $\eta=2g_{0}/\omega_{m}$ is used to characterize the optomechanical coupling [12], $\kappa$ is the cavity photon-number (energy) decay rate (linewidth); the field amplitude decays at $\kappa/2$ and $\Delta=\omega_{L}-\omega_{c}$ is the detuning between the laser frequency $\omega_{L}$ and the bare cavity resonance $\omega_{c}$. The mechanical damping rate parameter is denoted by $\gamma$. Note that we denoted the mechanical frequency by $\Omega_{m}$ in the first part of the paper and we now use $\omega_{m}$.
The Kerr effect causes a nonlinear, intensity-dependent shift in resonance frequency that is straightforward for the mean field but more complex for fluctuations due to quantum effects. The difference between these two effective detunings arises because the mean field experiences a static shift proportional to mean photon number, while fluctuations experience additional dynamical corrections reflecting their nonclassical nature.
The term effective detuning $\Delta_{\rm eff}$ refers to the fact that the cavity detuning is dynamically modified due to the presence of Kerr nonlinearity and depends on the intracavity amplitude. This means that the resonance of the cavity is shifted depending on the photon number inside the cavity, which leads to a nonlinear self-consistency for the detuning. When analyzing the system at the level of the mean field (the coherent amplitude $\alpha$), the Kerr nonlinearity induces a shift in the cavity resonance proportional to the mean photon number, so the effective detuning for the mean field is [12]
$\Delta_{\rm eff}=\Delta+2K\sum_{n}\left|\alpha_{n}\right|^{2},$
where $K$ denotes the strength of an effective Kerr nonlinearity with $K=\frac{g_{0}^{2}}{\omega_{m}}$ and $|\alpha|^{2}$ is the mean photon number. This accounts for the shift seen by the average field in the cavity.
For the quantum fluctuations around the mean field, the situation is more subtle. Fluctuations “see” a slightly different detuning because their dynamics are affected by both the mean field shift and the changing photon number, which can be time-dependent and has its own variance. The expression for the effective detuning for fluctuations must take into account higher-order corrections and the time dependence in the Kerr interaction. This leads to
$\tilde{\Delta}_{\rm eff}=\Delta+4K\sum_{n}\left|\alpha_{n}\right|^{2}.$
The distinction between the effective detuning for the mean field and that for the fluctuations arises from the different ways the Kerr nonlinearity affects these two components of the optical field. The difference can often be neglected if the fluctuations are small compared to the mean field. This is typically the case in the regime where the system is well away from bistability or instability thresholds, and photon number fluctuations are much smaller than the mean photon number (i.e., coherent state or weak nonlinearity). These two expressions for the effective detunings are derived in [12] with their semipolaron formalism.
The static approximation for the detuning is valid for a negligible Kerr parameter $K$ or in the weak driving limit. Indeed, in the $\left|\alpha(t)\right|^{2}\ll 1$ limit the bare detuning is just shifted by the constant Kerr term and there is no need to distinguish between $\Delta_{\rm eff}$ and $\widetilde{\Delta}_{\rm eff}$. The resulting Kerr shift of the detuning $\Delta_{\rm eff}=\Delta+K$ in that limit is also derived from the semipolaron transformation in [12], not as a phenomenological correction. It comes from how the radiation pressure interaction effectively displaces the mechanical oscillator based on the cavity photon number. This displacement, when properly accounted for, feeds back into the cavity frequency, leading to that renormalized detuning. This reflects a self-consistent frequency shift of the cavity due to the presence of photons — the hallmark of the optical Kerr effect, here induced via radiation pressure. Previous work [16] had noticed the shift in numerical studies and in [25] it had been noted that a renormalized detuning is essential for capturing nonclassical limit-cycle behavior.
@@
A consideration of both the static and the dynamical setting is alternatively presented throughout the whole discussion in Ref. [12]. Here, we first discuss the static case, valid for $\left|\alpha(t)\right|^{2}\ll 1$ and then in Section B.1 below, we discuss the dynamical case by evaluating $\Delta_{\rm eff}(r)$ in full, in contrast to the sideband-resolved approximation of [12]. Indeed, these sums have been evaluated in a close-to-resonance approximation [12], where one discards an infinite number of terms. Using the summation formulas of appendix B, we can now provide exact expressions for the complete resummation. For the drift, we obtain
@@
Details of the calculation are given in appendix B and below we will discuss its asymptotic value for large $\eta r$. For the diffusion coefficient, we find
@@
[…]
@@
$\tilde{\nu}=\frac{\tilde{\Delta}_{\rm eff}}{\omega_{m}}+\frac{i\kappa}{\omega_{m}}$ and the constants $B$, $C_{1}$, $C_{2}$, $C_{3}$ are given in appendix B.
@@ III.1.1 Diffusion in the Wigner case
The Fokker-Planck equation (25) can also be written for the Wigner function, instead of for $Q$ [12]. It has the same drift coefficient, and a diffusion given by (see (B.20) in [12])
@@
All the arguments of the Bessel functions are $\eta r$. This expression plays a relevant role in the dynamics of limit cycles. The infinite sums can again be computed exactly by using the summation formulas of Appendix B. The final result is a long expression and we will not need it here, but $D_{W}$ drastically simplifies when $\tilde{\Delta}_{\rm eff}=\Delta_{\rm eff}$. This identification is justified in the sideband-resolved regime studied in [12]. We obtain (see Appendix B for details)
@@
In figs. 3, 4, we compare our exact formula for $D_{W}$ with $D_{W}^{(0)}$ for $\kappa=0.1$, $\omega_{m}$ (resolved sideband) and two different values of $\Delta_{\rm eff}$. We see that the approximation (36) is very inaccurate for $\Delta_{\rm eff}=0.6\,\omega_{m}$, with a difference that oscillates between 20 and 75 percent (see fig. 4). For $\Delta_{\rm eff}=0.1\,\omega_{m}$, the difference can reach 20 per cent for some values of $r$.
@@
Figure 3: Optical part of the Wigner diffusion from approximation (36) (blue) and for the exact formula (35) (red) for $\kappa=0.1\omega_{m}$. a) $\Delta_{\rm eff}=0.1\,\omega_{m}$. The approximation (36) reproduces with high accuracy the exact curve. b) $\Delta_{\rm eff}=0.6\,\omega_{m}$. The approximation (36) is no longer applicable as higher $n$ terms in the infinite sum (34) become important.
Figure 4: Fractional differences as a function of $r$. For $\Delta_{\rm eff}=0.1\,\omega_{m}$ (brown curve) the difference reaches up to 20 per cent, whereas for $\Delta_{\rm eff}=0.6\,\omega_{m}$ (black curve) the difference is over 70 per cent for some values of $r$.
III.2 Asymptotic formulas, classical optomechanics cycles and the dynamical setting.
@@
To simplify the notation, here we write $\Delta\equiv\Delta_{\rm eff}(r)$. This formula exhibits the periods of oscillations in $r$ and in $\Delta/\omega_{m}$. A surprising effect predicted by the asymptotic formula for $\mu(r)$ is that the second term in (38) identically vanishes for
@@
Thus, for these special values of $\Delta$, $\mu(r)=-\gamma r+O(1/r)$. Oscillations are suppressed by a factor $1/r$ originating from subleading terms in the asymptotic expansion of the Bessel functions. This effect is shown in fig. 5 and is a result of the full resummation. It is not seen if one truncates the sum to few values of $n$, as in a close-to-resonance approximation.
An important challenge is to observe this effect experimentally. This requires designing the cavity such that the detuning, $\Delta=\omega_{L}-\omega_{c}$, equals an integer multiple of the mechanical frequency, $\omega_{m}$. In this case, measuring the drift $\mu(r)$ as a function of amplitude should reveal the behavior shown in Fig. 5c: the product $r^{2}\mu$ oscillates, with its maxima approaching a constant value for large amplitudes $r$. This behavior is clearly distinguishable from the case $\Delta\neq n\omega_{m}$, where the maxima of $r^{2}\mu$ increase linearly with the amplitude.
@@ III.2.2 Dynamical setting
The infinite sum can again be computed using the summation formulas of Appendix B by noting that this can be re-writen as
@@
This can be solved numerically, or analytically when $\Delta_{\rm eff}\ll\omega_{m}$. In this case
@@
This formula encodes the information of all secondary resonances. Notice that (46) is also valid in the opposite regime of the unresolved sideband (bad cavity limit) and therefore it provides a good analytical evaluation for $\Delta_{\rm eff}$ for $c=\kappa/2\omega_{m}>1$ too, a regime not covered in [12]. This regime is of great physical interest because it is easily achieved as the mechanical oscillator size increases, since larger oscillators are characterised by lower frequencies.
It is instructive to compare the $\Delta_{\rm eff}$ obtained from the exact summation formula with that derived from an approximation based on retaining the main resonance. While the difference depends on the specific parameter values, it is possible to anticipate the conditions under which this discrepancy becomes significant. If $\kappa\ll\omega_{m}$ and $\Delta/\omega_{m}$ is very close to a resonance, i.e. very close to an integer $\Delta\approx n_{0}\omega_{m}$, then the term $n=n_{0}$ dominates over all other terms. Any other $n$-term in the sum is suppressed by a factor $\frac{\kappa^{2}}{\omega_{m}^{2}}\frac{1}{(n-n_{0})^{2}}\ll 1$. In this case, retaining the main resonance provides a good approximation.
However, the most important deviations between the exact and the approximate formula arise when $\Delta_{\rm eff}$ is not close to a resonance. Let us consider an example. If $KE^{2}$ is small, one has $\Delta_{\rm eff}\sim\Delta$ in a leading-order approximation. If $\Delta/\omega_{m}$ is a half integer, then oscillations in the amplitude $r$ are suppressed in the exact formula, as can be seen from (46) by virtue of the fact that $\cos(\pi\frac{\Delta_{\rm eff}}{\omega_{m}})=0$. However, oscillations are not suppressed in a truncation that keeps only one or two terms of the sum. This generally leads to a significative difference in the numerical results. The difference can also be especially relevant, in particular, when it implies a change of sign in the total $\Delta_{\rm eff}$ with regards to the sign of $\Delta$ [25, 12] (recall the discussion of the static approximation in Sec III.A).
In addition to these numerical differences, it is important to note that an approximation based on a truncated sum fails to capture certain properties that emerge only upon full resummation over $n$ – specifically, those arising at special values of $r$ and $\Delta_{\rm eff}$, for which the trigonometric functions in (45) may vanish.
@@ IV Outlook
The technical advance of this paper lies in the formulation of a compact, simple formula that comprehensively encompasses all terms within the summation. This provides an analytical understanding of the parameter space and offers insight into the underlying physics of the various processes. Additionally, it facilitates the application of asymptotic expressions, resulting in expressions that are expressed in terms of uniquely elementary functions. Since these are simple trigonometric functions, further elucidation, for example of the zeros [10], becomes a particularly straightforward task, helping to identify special points in regions of the parameter space in limits where a resolved sideband approximation, which retains only a few resonant terms in the summation, does not reach. There are also open problems already in the context of the present discussion, for example, whether the more complete analysis of Wigner diffusion presented here also allows a more detailed analysis of energy fluctuations by evaluating the Fano factor, as in [12], but with full summation. All these considerations require no further mathematical developments than those presented here or in the mathematically related works [9, 10].
@@ Acknowledgements.
We thank Professors David Pérez-García, Oriol Romero-Isart, Patrick Maurer, Wenjie Ma and Jiteng Sheng for discussions and/or correspondence at different stages of this work. The work of MT was started at Universidad Complutense de Madrid (UCM) and continued and finished after moving to the Shanghai Institute for Mathematics and Interdisciplinary Sciences (SIMIS). This work was supported by the Shanghai Institute for Mathematics and Interdisciplinary Sciences (SIMIS) under the Applied and Interdisciplinary Collaboration Project. M.T. gratefully acknowledges the resources and facilities provided by SIMIS, which were essential for the completion of this research. JGR acknowledges financial support from grant 2021-SGR-249 (Generalitat de Catalunya) and by the Spanish MCIN/AEI/10.13039/501100011033 grant PID2022-126224NB-C21. MT acknowledges previous financial support from FEI-EU-22-06, funded by Universidad Complutense de Madrid, and grant PID2020-113523GB-I00, funded by the Spanish Ministry of Science and Innovation.
Appendix A Fokker-Planck description of limit cycles
The paper [12] develops an analytical Fokker-Planck formalism for describing mechanical limit cycles in the quantum regime of optomechanical systems. After summarizing the conventions and notation used through the main text, we present a summary of this formalism.
Conventions and notation
Drive and energy. We reserve $E$ only for the laser drive rate (units of s-1), sometimes also called drive amplitude, related to the input field by $E=\sqrt{\kappa_{\rm ex}}\,\alpha_{\rm in}$ and $P_{\rm in}=\hbar\omega_{L}|\alpha_{\rm in}|^{2}$. This is not to be confused with the mechanical energy, denoted $\mathcal{E}=\hbar\omega_{m}r^{2}$. We use $K=g_{0}^{2}/\omega_{m}$ and $\eta=2g_{0}/\omega_{m}$.
Detuning and effective detunings. $\Delta=\omega_{L}-\omega_{c}$. We use $\Delta_{\rm eff}=\Delta+2K\sum_{n}|\alpha_{n}|^{2}$ for the optical mean field and $\tilde{\Delta}_{\rm eff}=\Delta+4K\sum_{n}|\alpha_{n}|^{2}$ for fluctuations (twice the shift). In the sideband-resolved regime studied in [12], these coincide to leading order.
Damping-rate conventions (factor-of-two). There are two conventions as explained in [12] (i) The one used in [12] and hence in our second part: $\kappa$ denotes the field-amplitude decay rate in the Heisenberg–Langevin equation, $\dot{a}=(i\Delta-\kappa)a+\sqrt{2\kappa}\,a_{\rm in}$; the photon-number (energy) decay rate is $2\kappa$. (ii) The one in the earlier works [7, 11] (and hence used in our Part I): $\kappa_{\rm(energy)}$ denotes the photon-number decay rate and the field amplitude decays at $\kappa_{\rm(energy)}/2$. Therefore a factor of 2 between the two. We adhere to these respective conventions within each part.
Starting Point: Optomechanical Master Equation
The dynamics of the coupled cavity-mechanical system is governed by the standard quantum master equation [3]:
$\frac{d}{dt}\rho=(\mathcal{L}_{m}+\mathcal{L}_{c}+\mathcal{L}_{\text{int}})\rho,$
where:
  • $\mathcal{L}_{m}$: Liouvillian for the mechanical oscillator,
  • $\mathcal{L}_{c}$: Liouvillian for the cavity,
  • $\mathcal{L}_{\text{int}}$: Cavity-mechanics interaction term.
Explicitly, these are:
$\mathcal{L}_{m}\rho$ $=-i[\omega_{m}b^{\dagger}b,\rho]+\gamma(\bar{n}+1)\mathcal{D}[b]\rho+\gamma\bar{n}\mathcal{D}[b^{\dagger}]\rho,$
$\mathcal{L}_{c}\rho$ $=-i[-\Delta a^{\dagger}a-iE(a-a^{\dagger}),\rho]+\kappa\mathcal{D}[a]\rho,$
$\mathcal{L}_{\text{int}}\rho$ $=-i[-g_{0}a^{\dagger}a(b+b^{\dagger}),\rho],$
where $a,b$ are photon and phonon (mechanical) annihilation operators, respectively, and $g_{0}$ is the single-photon optomechanical coupling. The frequency of the mechanical oscillator is $\omega_{m}$, its amplitude damping rate is $\gamma=\frac{\omega_{m}}{Q_{m}}$, and the mean phonon number in thermal equilibrium is $\overline{n}$. The Lindblad operators are denoted by $\mathcal{D}[a]$ [12] and $\kappa$ is the cavity photon-number (energy) decay rate (linewidth); the field amplitude decays at $\kappa/2$ and $\Delta$ is the detuning from cavity resonance of the driving field $E$.
2. Semipolaron Transformation and Phase-Space Representation
The formalism uses a phase-space representation (such as the Husimi Q-function) for the mechanical oscillator, introducing a joint phase-space operator $\sigma(\beta,\beta^{*})$. Through a “semipolaron” transformation the optical Kerr nonlinearity is separated and the nonlinearity of the optomechanical interaction can be considered in an analytically tractable way. This transformation is designed to facilitate the analysis of nonlinear quantum dynamics in cavity optomechanical systems, especially when studying limit cycles beyond simple linearized theories.
This results in a hybrid description where the mechanical oscillator is described in phase space while the optical cavity remains in Hilbert space. The main feature of the semipolaron transformation is to partially “dress” the cavity field, introducing a phase shift that depends on the mechanical oscillator’s position in phase space. The transformation is formally defined as:
$\widetilde{\sigma}(\beta,\beta^{*},t)=e^{\eta(\beta-\beta^{*})a^{\dagger}a/2}\sigma(\beta,\beta^{*},t)e^{-\eta(\beta-\beta^{*})a^{\dagger}a/2}\ ,$
or equivalently as a unitary transformation:
$\widetilde{\sigma}(\beta,\beta^{*},t)=e^{i\theta(\beta,\beta^{*})a^{\dagger}a}\sigma(\beta,\beta^{*},t)e^{-i\theta(\beta,\beta^{*})a^{\dagger}a}\ ,$
with $\theta(\beta,\beta^{*})=\eta\operatorname{Im}(\beta)$ and $\eta=2g_{0}/\omega_{m}$. This approach explicitly separates the cavity’s Kerr nonlinearity (arising from the radiation pressure coupling) from the optomechanical interaction, while still retaining a residual interaction term. Importantly, it preserves a clear interpretation of the mechanical state by allowing for a well-defined reduced density operator for the mechanics.
This transformation is named “semipolaron” because it bears similarity to the full polaron transformation (which entirely diagonalizes the interaction at the cost of entangling cavity and mechanics), but only “partially” transforms: it keeps the two subsystems largely separate, allowing calculations in which second-order perturbation theory can be applied to the remaining interaction term.
In contrast, the standard polaron transformation would remove the optomechanical interaction completely but prevent straightforward access to the reduced state for the mechanical oscillator, which is essential when studying quantum limit cycles. The Liouvillian for the cavity acquires an effective Kerr nonlinearity, and the drive becomes phase-modulated in a way that depends on the mechanical oscillator.
3. Adiabatic Elimination and Effective Fokker-Planck Equation
Now the jump operator for cavity dissipation remains unchanged—unlike in a full polaron picture—while the crucial relation between $\widetilde{\sigma}(\beta,\beta^{*},t)$ and the mechanical Q-function is preserved. This separation enabled the authors in [12] to perform an adiabatic elimination of the cavity. By adiabatically eliminating the cavity and focusing on the mechanical degree of freedom, the effective Fokker-Planck equation (FPE) for the reduced Q-function (or equivalently, the Wigner function, with suitable corrections) of the mechanical oscillator is obtained. Schematically,
$\frac{\partial}{\partial t}Q(\beta,\beta^{*})=\text{Drift}+\text{Diffusion}+\text{Coupling terms}.$
After phase averaging and in polar coordinates ($\beta=re^{i\varphi}$), the FPE reduces to a 1-dimensional form for the amplitude $r$:
$\frac{\partial}{\partial t}Q(r)=-\partial_{r}[\mu(r)Q(r)]+\partial_{r}^{2}[D(r)Q(r)],$
where the key quantities (corresponding to (26) and (27) in the main text) are:
Drift coefficient:
$\mu(r)=-\gamma r+\sum_{n}g_{0}E^{2}\,\mathrm{Im}\left[\frac{J_{n-1}(\eta r)J_{n}(\eta r)}{h_{n-1}h^{*}_{n}}\right]$
Diffusion coefficient:
$D(r)=\frac{\gamma(\bar{n}+1)}{2}+\sum_{n}\frac{g_{0}^{2}E^{2}}{2}\left[\frac{\kappa J_{n}^{2}(\eta r)}{|h_{n}|^{2}|\tilde{h}_{n-1}|^{2}}-\mathrm{Re}\left(\frac{J_{n-2}(\eta r)J_{n}(\eta r)}{\tilde{h}_{n-1}h^{*}_{n-2}h_{n}}\right)\right]$
with
  • $\kappa$: cavity photon-number (energy) decay rate (linewidth); the field amplitude decays at $\kappa/2$,
  • $J_{n}$: Bessel functions (from sidebands generated by mechanical motion),
  • $h_{n}=\kappa+i(n\omega_{m}-\Delta_{\text{eff}})$,
  • $E$: laser drive rate,
  • $\Delta_{\text{eff}}$: effective detuning including Kerr effect.
This formalism captures nonlinear limit-cycle dynamics in the quantum regime. This Fokker-Planck approach bridges optomechanical laser theory with quantum statistical mechanics, enabling tractable yet powerful predictions for quantum effects in experimentally relevant parameter regimes. In this work we have evaluated analytically both the drift and diffusion coefficient and the ensuing asymptotic behavior in different regimes and its consequences, thereby extending the analytical results of the Fokker-Planck formalism of [12].
Appendix B Bessel summations and diffusion coefficients